solve for a in terms of b when 2b=3(a-4)
Answer by umail08(87)

2b = 3(a-4)
2b = 3a – 12
2b + 12 = 3a – 12 + 12
2b + 12 = 3a … divide both sides by 3
2/3b + 4 = a